2. Measuring standalone risk using realized (historical) data Aa Aa Returns earned over a given time period are called realized returns. Historical data on realized returns is often used to estimate future results. Analysts across companies use realized stock returns to estimate the risk of a stock. Consider the case of Falcon Freight Inc. (FF): Five years of realized returns for FF are given in the following table. Remember: 1. While FF was started 40 years ago, its common stock has been publicly traded for the past 25 years 2. The returns on its equity are calculated as arithmetic returns. 3. The historical returns for FF for 2012 to 2016 are: 2012 2013 2014 2015 2016 Stock return 25.00% 17.00% 30.00% 42.00% 13.00% Given the preceding data, the average realized return on FFs stock is The preceding data series represents standard deviation of FFs historical returns is of FFs historical returns. Based on this conclusion, the If investors expect the average realized return from 2012 to 2016 on FFs stock to continue into the future, its coefficient of variation (CV) will be
Answer #1

Part1:
The average realized return is (25+17+30+42+13)/5=25.4%
Part 2:
Preceding data series represents a sample.
Part 3:
To find sample standard deviation we will follow the following steps:
We have calculated the average or mean return as 25.4%
Returns given in the observations are:
25%, 17%, 30%, 42%, 13%
Now, we will subtract the average value from these observation values and square the results.
(25-25.4)^2=.16
(17-25.4)^2=70.56
(30-25.4)^2=21.16
(42-25.4)^2=275.56
(13-25.4)^2=153.76

Now, we need to add the values or take summation of the values.
.16+70.56+21.16+275.56+153.76=521.2
Next step is to find the value of n-1 where n is the number of sample observations.
Here n=5 so n-1=4
Now, we need to divide the value 4 with 521.2, that will give 521.2/4=130.3
To find standard deviation, we need to take square root of 130.3 which is equal to 11.4149.
Thus, standard deviation is 11.4 (approximately)

Part 4:
Coefficient of variation is the ratio of standard deviation/average
Coefficient of variation=11.4149/25.4=.4494=0.45(approximately)